OpenStudy (anonymous):

If you wanted to eliminate the y variable by adding the two equations in the following system, what could be your first step? 3x + y = 6 4x − 5y = 10 Multiply the first equation by 5. Multiply the first equation by −5. Multiply the first equation by −4. Add the two equations together.

OpenStudy (swissgirl):

Well what happens when u multiply the first equation by 5?

OpenStudy (anonymous):

i dunno is the first one the answer

OpenStudy (swissgirl):

well just try it out multiply each term by 5 5(3x+y=6)

OpenStudy (anonymous):

15 + y=6

OpenStudy (swissgirl):

well we use distributive property and multiply every term 15x+5y=30

OpenStudy (swissgirl):

now add the 2 equation 15x+5y=30 4x-5y=10 ------------ 19x +0y=40

OpenStudy (swissgirl):

as u see there is no y left. There are 0 y's meaning that we have eliminated the y

OpenStudy (anonymous):

so the answer is the first one right

OpenStudy (swissgirl):

yuppppp ditty dup

OpenStudy (swissgirl):

Awesome Job here is some dog food

OpenStudy (anonymous):

-_-